1) Suppose the real rate is 3.45 percent and the inflation rate is 2.2 percent. What rate would you expect to earn on a Treasury bill?

A) 1.25 percent

B) 7.75 percent

C) 5.73 percent

D) 6.56 percent

E) 3.30 percent


2) The common stock of Dayton Repair sells for $ 47.92 a share. The stock is expected to pay $ 2.28 per share next year when the annual dividend is distributed. The company increases its dividends by 1.65 percent annually. What is the market rate of return on this stock?

A) 6.14 percent

B) 6.41 percent

C) 9.92 percent

D) 4.84 percent

E) 7.28 percent


3) Farmco just paid its annual dividend of $ .32 per share. The dividends are expected to grow at 25 percent annually for the next 4 years and then level off to an annual growth more of 3 percent indefinitely. What is the price of this stock today given a required return of 15 percent?

A) $ 9.01

B) S 6.03

C) $ 7.54

D) $ 5.42

E) $ 7.32


4) At 5 percent interest, how long would it take to triple your money?

A) 26.55 years

B) 25.64 years:

C) 24.87 years

D) 22.52 years

E) 20.01 years
Answer #1

1). Nominal Rate = [(1 + Real Rate) * (1 + Inflation)] - 1

= [(1 + 0.0345) * (1 + 0.022)] - 1 = 1.0573 - 1 = 0.0573, or 5.73%

Hence, Option "C" is correct.

2). r = [D1 / P0] + g

= [$2.28 / $47.92] + 0.0165 = 0.0476 + 0.0165 = 0.0641, or 6.41%

Hence, Option "B" is correct.

3). P0 = [{D0 * (1 + g1)} / (1 + r)] + [{D0 * (1 + g1)2} / (1 + r)2] + [{D0 * (1 + g1)3} / (1 + r)3] + [{D0 * (1 + g1)4} / (1 + r)4] + [{D0 * (1 + g1)4 * (1 + gC)} / {(r - gC) * (1 + r)4}]

= [{$0.32 * (1 + 0.25)} / (1 + 0.15)] + [{$0.32 * (1 + 0.25)2} / (1 + 0.15)2] + [{$0.32 * (1 + 0.25)3} / (1 + 0.15)3] + [{$0.32 * (1 + 0.25)4} / (1 + 0.15)4] + [{$0.32 * (1 + 0.25)4 * (1 + 0.03)} / {(0.15 - 0.03) * (1 + 0.15)4}]

= $0.35 + $0.38 + $0.41 + $0.45 + $3.83 = $5.42

Hence, Option "D" is correct.

4). n = [log(FV/PV)] / [log(1 + r)]

= [log(3)] / [log(1 + 0.05)]

= 1.0986 / 0.0488 = 22.52 years

Hence, Option "D" is correct.